OpenStudy (anonymous):

you roll a number cube twice. find P(even, then not 2).

OpenStudy (ddcamp):

Assuming a six sided die: P(even) = 1/2 P(not 2) = 5/6 P(both) = 5/12

OpenStudy (anonymous):

thx